Your grandfather invested $1,000 in a stock 36 years ago. Currently, the value of his account is $318,000. What is his geometric return over this period? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1
Investment                                   1,000
Current value                               318,000
Time                                         36
318000= 1000*(1+r)^36
318000/1000= (1+r)^36
318 (1+r)^36
318^(1/36)= 1+r
1.173577743 1+r
r= 0.173577743
r= 17.36%